\begin{thebibliography}{10}

%\bibitem{AbK} M. Abouzaid, T. Kragh: Simple homotopy equivalence of nearby Lagrangians, {\it Acta Math.} {\bf 220} (2018), no. 2, 207-237.

\bibitem{A1} H. Alexander: Gromov’s method and Bennequin’s problem, {\it Invent. math.} {\bf 125} (1996), 135-148.

\bibitem{A2} H. Alexander: Disks with boundaries in totally real and Lagrangian manifolds, {\it Duke Math. J.} {\bf 100} (1999), No. 1, 131-138.

\bibitem{AHS}  M. F. Atiyah, N. J. Hitchin, I. M. Singer: Self-duality in four-dimensional Riemannian geometry, {\it Proc. Roy. Soc. London Ser. A} {\bf  362} (1978), no. 1711, 425-461.

%\bibitem{Au} M. Aubin: Fibres normaux d'immersions en dimension double, points doubles d'immersions iagrangiennes et plongements totalement r\'eels, {\it  Comment. Math. Helvetici} {\bf 63} (1988),  593-623. 

\bibitem{CGT} J. Cheeger, M. Gromov, M. Taylor: Finite propagation speed, kernel estimates for functions of the Laplace operator, and the geometry of complete Riemannian manifolds, {\it J. Differential Geom.} {\bf 17} (1982), 15-53.

\bibitem{ChenFraser1} J. Chen, A. Fraser, C. Pang: Minimal immersions of compact bordered Riemann surfaces with free boundary, {\it Trans. Amer. Math. Soc.} {\bf 367} (2015), no. 4, 2487-2507.

\bibitem{ChenFraser2} J. Chen, A. Fraser: Holomorphic Variations of Minimal Disks with Boundary on a Lagrangian Surface, {\it Canad. J. Math.}, Vol. {\bf 62} (2010), 1264-1275.

\bibitem{C-G} K. Cieliebak, E. Goldstein: A note on the mean curvature, Maslov class and symplectic area of Lagrangian immersions, {\it J. Symplectic Geom.} {\bf 2} (2004), no. 2, 261-266.

\bibitem{Da} P. Dazord: Sur la g\'eome\'etrie des sous-fibr\'es et des feuilelletages lagrangiens, {\it Ann. Sci. \'Ecole Norm. Sup.}, {\bf 14} (1981), 464-480.

\bibitem{Du} J. Duval: Convexit\'e rationnelle des surfaces Lagrangiennes, {\it Invent. math.} {\bf 104} (1991), 581-599.

\bibitem{DS1} J. Duval, N. Sibony:  Polynomial convexity, rational convexity, and currents, {\it Duke Math. J.} {\bf 79} (1995), 487-513.

\bibitem{E-L} J. Eells, L. Lemaire: Selected topics in harmonic maps, CBMS Regional Conference Series in Mathematics, 50. Published for the Conference Board of the Mathematical Sciences, Washington, DC; by the American Mathematical Society, Providence, RI, 1983..

\bibitem{Gruter} M. Gr\"uter: Regularity at the free boundary of two-dimensional stationary harmonic maps, {\it Ann. Inst. H. Poincar\'e C Anal. Non Lin\'eaire} {\bf 15} (1998), no. 2, 151-167. 

\bibitem{Gromov} M. Gromov: Pseudo holomorphic curves in symplectic manifolds, {\it Invent. math.} {\bf 82} (1985), 307-347.

%\bibitem{GH} P. Griffiths, J. Harris: Principles of Algebraic Geometry,  Wiley Classics Library. John Wiley \& Sons, Inc., New York, 1994.

\bibitem{KM} J. L. Koszul, B. Malgrange: Sur certaines fibrees complexes, {\it Arch. Math.} (Besel) {\bf 9} (1958), 102-109.

\bibitem{Jost1} J. Jost: Continuity of minimal surfaces with piecewise smooth free boundaries, {\it Math. Ann.} {\bf 276} (1987), no. 4, 599-614.

\bibitem{Li} A. Lichnerowicz: Applications harmoniques et vari\'et\'es k\"ahl\'erienes, {\it Symp. Math.} III, Bologna, (1970), 341-402. 

\bibitem{MS} D. McDuff, D. Salamon: $J$-holomorphic Curves and Symplectic Topology, Amer. Math. Soc. Colloquium Publications, {\bf 52}, {\it American Mathematical Soiety, Providence, RI}, 2012. 

\bibitem{MSY} W. Meeks, L. Simon, S. T. Yau: Embedded minimal surfaces, exotic spheres, and manifolds with positive
Ricci curvature, {\it Ann. of Math.} {\bf 116} (1982), 621-659.

\bibitem{MM} M. Micallef, J. D. Moore: Minimal two-spheres and the topology of manifolds with positive curvature on totally isotropic two-planes, {\it Ann. of Math.} (2) {\bf 127} (1988), no. 1, 199-227.

\bibitem{Miu} T. Miura:  A diameter bound for compact surfaces and the Plateau-Douglas problem, {\it Ann. Sc. Norm. Super. Pisa Cl. Sci.} (5) {\bf 23} (2022), no. 4, 1707-1721.

\bibitem{Mor} J. M. Morvan: Quelques invariants topologiques en g\'eom\'etrie  symplectique, {\it Ann. Inst. H. Poincar\'e, A} {\bf 38} (1983), 349-370.

\bibitem{Mo} C. B. Morrey Jr.: The problem of Plateau on a Riemannian manifold, {\it Ann. of Math.}
 {\bf 49} (1948), 807-851. 

\bibitem{Ono} H. Ono: Integral formula of Maslov index and its applications, {\it Japaness J. Math. (N.S.)}, Vol {\bf 30}, No. 2 (2004), 413-421. 

\bibitem{Pacini} T. Pacini: Maslov, Chern-Weil and mean curvature, {\it  J. Geom. Physics} {\bf 135} (2019), 129-134.

\bibitem{SU} J. Sacks, K. Uhlenbeck: The existence of minimal immersions of 2-spheres, {\it Ann. of Math.} 
 {\bf 113} (1981), no. 1, 1-24.

\bibitem{SaAg} A. Salimov, F. Agca: Some perperties of Sasakian metrics in contangent bundles, {\it Mediter. J. Math.} {\bf 8} (2011), 243-255.

\bibitem{Tond} P. Tondeur: Structure presque K\"ahl\'erienne naturelle sur le fibr\'e des vecteurs covariants d'une vari\'et\'e riemanienne, {\it C.R. Acad. Sci. Paris} {\bf 254} (1962), 407-408.

%\bibitem{V} I. Vaisman: Symplectic Geometry and Secondary Characteristic Classes, Progress in Mathematics, Vol {\bf 72}, Springer, 1987. 

\bibitem{Vianna} R. Vianna: On exotic Lagrangian tori in $\C{\mathbb P^2}$, {\it Geom. Topol.} {\bf 18} (2014), no. 4, 2419-2476.

\bibitem{Ye} R. Ye: On the existence of area-minimizing surfaces with free boundary, {\it Math. Z.} {\bf 206} (1991), no. 3, 321-331.
\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amsmath}
\usepackage{graphicx}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage[utf8]{inputenc}
\usepackage{comment}
\usepackage{xcolor}
\usepackage{lipsum}
\usepackage{ifthen}
\usepackage{stmaryrd}

\usepackage{tikz}

\usepackage{newtx}
\usepackage{relsize}

\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}

\theoremstyle{remark}
\newtheorem{rem}[thm]{Remark}
\numberwithin{equation}{section}

\newcommand{\R}{\mathbb R}

\newcommand{\A}{\mathcal{A}}

\newcommand{\ed}{\end {document}}

\def\C{{\mathbb C}}
\def\S{{\mathbb S}}
\def\p{\partial}

\def\vp{\varphi}

\def\vp{\varphi}

\def\n{\nabla}

\def\E{{\bf E}}
\def\A{\mathcal A}

\def\C{{\mathbb C}}
\def\S{{\mathbb S}}
\def\p{\partial}

\def\vp{\varphi}

\def\vol{\mbox{vol}}

\begin{document}

\title[Holomorphic disks]{Holomorphic disks with boundary on compact Lagrangian surface}
\author{Jingyi Chen}
\date{\today}
\thanks{2000 {\em Mathematics Subject Classification.} Primary 58E12; Secondary 53D12, 53C21. }
\thanks{The research is partially supported by NSERC Discovery Grant GR010074.}
\address{Department of Mathematics, The University of British
Columbia, Vancouver, B.C. V6T 1Z2, Canada}
\email{jychen@math.ubc.ca}

\begin{abstract}
Let $L$ be a compact oriented Lagrangian surface in a K\"ahler surface endowed with a complete Riemannian metric (compatible with the symplectic structure and the complex structure) with bounded sectional curvatures and a positive lower bound on injectivity radius. We show that for every nontrivial class $[\gamma]$ of the fundamental group $\pi_1(L)$ such that $\gamma$ bounds a topological disk in $M$, there exists a holomorphic disk whose boundary belongs to $L$ and is freely homotopic to $\gamma$ on $L$. This answers a question of Bennequin on existence of $J$-holomorphic disks. Nonexistence of exact Lagrangian embeddings of certain surfaces is established in such K\"ahler surface if the fundamental form is exact. In the almost K\"ahler setting, especially, the cotangent bundles of compact manifolds, results on nonexistence of $J$-holomorphic disks and existence of minimizers of the partial energies in the sense of A. Lichnerowicz are obtained. 
\end{abstract}

\maketitle

\section{Introduction}

In his famous work \cite{Gromov} on pseudo-holomorphic curves, M. Gromov proves, among many other important things, the existence of a nonconstant holomorphic map $u:(D,\p D)\to (\C^n,L)$ for any closed Lagrangian submanifold $L$ in $\C^n$, and he points out that this partially solves a conjecture of D. Bennequin on existence of a holomorphic disk in $\C^2$ with boundary in a totally real torus $T^2$  \cite[Theorem, Remark, p.313]{Gromov}. 
Since then, results confirming the existence of holomorphic disks with boundary lying on a totally real submanifold are obtained in specific situations; however, a totally real torus in $\C^2$ bounding no analytic disc is constructed explicitly (\cite{A1, A2}, \cite{Du}, \cite{DS1}). 

In this paper, we first focus on Bennequin's question for the Lagrangian case in complex dimension two. Specifically, we consider existence of holomorphic disks with boundary lying in a compact Lagrangian surface $L$ in a K\"ahler surface $M$. We assume that the Riemannian manifold $M$ is {\it homogeneously regular}, that is,  $M$ is complete and its injectivity radius is bounded below by a positive constant and its sectional curvatures are bounded (\cite{Mo}, \cite{MSY}). This requirement is imposed for the existence of minimal disks with free boundary on $L$ when $M$ is noncompact. 

Our main existence result is that each nontrivial class $[\gamma]\in\pi_1(L)$, while $\gamma\subset L$ bounds a topological disk in $M$, is representable in the sense of free homotopy on $L$ by the boundary of a nonconstant holomorphic disk as a branched minimal immersion, and the disk meets $L$ orthogonally along its boundary. This orthogonality with the fact that $L$ is Lagrangian turns out to be crucial in proving constancy of certain holomorphic functions in our arguments. 

\begin{thm}\label{represent} 
Let $M$ be a K\"ahler surface whose underlying Riemannian structure is homogeneously regular. Let $L$ be a compact oriented Lagrangian submanifold of $M$. Then there exists a generating set $\{\gamma_j\}$ for $\ker i_*$, where $i_* : \pi_1(L)) \to \pi_1(M)$ is the homomorphism induced by the inclusion $i:L\hookrightarrow M$, such that each $\gamma_j$ is freely homotopic to the boundary of a holomorphic disk that meets $L$ orthogonally along its boundary. The boundary Maslov class of each of these disks is nonzero. 
\end{thm}

In \cite[2.3]{Gromov}, the existence of holomorphic curves is obtained by directly dealing with the $\bar\p$-equation and compactness of pseudo-holomorphic curves. In contrast, we begin with minimal surfaces with free boundary on the restraining surface $L$. The existence and regularity theory (of free boundary minimal surfaces) is based on a ``replacement procedure" in \cite{Ye} or on Sacks-Uhlenbeck's perturbed energy method in \cite{ChenFraser1}. The boundaries of these minimal (in fact, area minimizing) disks form a generating set of $\ker i_*$. From the first variation formula of area of bordered surfaces, the minimal disks are orthogonal to $L$ along their boundaries. This sheds light on deducing holomorphicity of these minimal disks as $L$ is Lagrangian. This is further ensured if there are sufficiently many holomorphic variation vector fields with real part lying in the the tangent bundle of $L$ along the boundary of the disk. A precise counting of these admissible variations involves the boundary Maslov index and is established in \cite{ChenFraser2} (essentially, it is a Riemann-Roch theorem for a bundle pair over a bordered Riemann surface). A technical point important for the current work is the criterion in \cite[Theorem 1.1 (1)]{ChenFraser2} detecting holomorphicity via linear independence of the admissible variations at a single point on the boundary of the disk. 

The method outlined above extends Gromov's theorem on existence of holomorphic disks to more general types of $L,M$ and in each nontrivial class of $[\gamma]\in\pi_1(L)$ in $\ker i_*$, when $\dim_\C M=2$. 
In particular, it demonstrates existence of multiple holomorphic disks, for example: If $M$ is simply connected, the long exact sequence of homotopy groups
\begin{equation}\label{exact sequence}
\cdots \to \pi_2(M) \to \pi_2(M,L)\overset{\partial}{\to}\pi_1(L)\overset{i_*}{\to} \pi_1(M) \to\cdots
\end{equation} 
asserts ${\mbox{Im} }\,\partial=\ker i_* = \pi_1(L)$. When $L$ is a Lagrangian torus $T^2$ in $\C^2$, there exist at least two nonconstant holomorphic disks with boundary in $T^2$ representing distinct nontrivial classes in $\pi_1(T^2)\cong \mathbb Z \times\mathbb Z$. Existence of multiple holomorphic disks have been studied via techniques involve Floer's theory, see for example \cite{Vianna} and the reference therein. 

\vspace{.1cm} 

A striking consequence of Gromov's existence theorem on holomorphic disk is: There does not exist any compact embedded exact Lagrangian submanifold in $\R^{2n}$ with the standard symplectic structure; furthermore, this leads to the discovery of an exotic symplectic structure on $\R^{2n}$ and to the fact that a torus $T^2$ cannot be embedded as an exact Lagrangian surface in $\C^2$, see \cite{Gromov}.

We shall explore the exact Lagrangian embeddings. First, we prove nonexistence of exact Lagrangian embeddings in the K\"ahler setting: 

\begin{thm}\label{no exact} 
Let $M$ be a K\"ahler surface with exact fundamental form $\omega$ and compatible homogeneously regular metric. Let $L$ be a compact orientable surface embedded in $M$. If there is a loop $\gamma\subset L$ which is nontrivial in $\pi_1(L)$ but trivial in $\pi_1(M)$, then $L$ cannot be exact Lagrangian. 
\end{thm}

The above result is a consequence of Theorem \ref{represent} and Lemma \ref{homotopy}. The latter is an extension of Lichnerowicz's homotopy lemma for smooth maps between almost K\"ahler manifolds in terms of the partial energies $E'$ and $E''$ \cite{Li}, to the case that the domain has nonempty boundary with image in a Lagrangian submanifold under maps inside the relative homotopy class. 

\vspace{.1cm}

Next, we work with the almost K\"ahler setting: $M$ is equipped with a symplectic form $\omega$, an almost complex structure $J$ and a Riemannian metric $h$ which satisfy $\omega(X,Y)=h(JX,Y)$ and $h(JX,JY)=h(X,Y)$ for all $X,Y\in T_pM$ and all $p\in M$. An almost K\"ahler structure is K\"ahler if $\n J=0$. 
When $M$ is almost K\"ahlerian, the interplay between existence of holomorphic disks and existence of exact Lagrangian embeddings 
bocomes delicate via our approach. A reason is that it is unclear how to determine the holomorphicity of an area minimizing disk when the condition $\n J=0$ is dropped , even for the cotangent bundle of a compact orientable surface (other than a torus). 

\begin{thm}\label{not exist} 
Let $M$ be a manifold with an almost K\"ahler structure $(\omega, J, g)$ and $\omega$ is exact. Let $L$ be an exact Lagrangian submanifold in $M$. Then 
\begin{enumerate}
\item There is no nonconstant $J$-holomorphic map $u:(D,\p D)\to (M,L)$.
\item If $g$ is homogeneously regular, then there is a generating set $\{\gamma_j\}$ of $\ker i_*$ where $i_*:\pi_1(L)\to\pi_1(M)$ is induced by $i:L\hookrightarrow M$ such that each $\gamma_j$ is freely homotopic to the boundary of an energy minimizing harmonic map $u:(D,\p D)\to(M,L)$ which minimizes both $E',E''$ in the relative homotopy class. 
\end{enumerate}
\end{thm}

Lastly, we focus on the cotangent bundle $T^*M$ of a compact manifold $M$, endowed with the canonical symplectic structure $\omega_{\tiny{\rm can}}$. 
For any Riemannian metric $g$ on $M$, the Sasakian metric $^{\tiny{{\mbox S}}}\!g$ on $T^*M$ is complete (see Proposition \ref{complete}). The sectional curvatures of $^{\tiny{\mbox S}}\!g$ are unbounded (unless $g$ is special, e.g. flat), while the unboundedness of the curvature is at most of a polynomial order along the fibres over $M$, see \cite{SaAg}. Choosing an appropriate conformal factor, there exists a homogeneously regular metric $g_{_f}=e^{2f}\,{^{\tiny \mbox{S}}\!g}$. The injectivity radius lower bound follows from a theorem of Cheeger, Gromov and Taylor in \cite{CGT}, after establishing a lower bound on volume of geodesic balls. 
The existence theory of free minimal disks then applies to $(T^*M,g_{_f})$. 

Although $\omega_{\rm can}$ and $g_{_f}$ determine an almost complex structure on $T^*M$, the triple hardly ever provides an almost K\"ahler structure on $T^*M$ due to the conformal factor. It turns out that by estimating the diameter of the minimal disk we can show that the minimal disk is in fact contained in a region $\Omega$ on which the conformal metric coincides with the Sasakian metric. Consequently, the almost K\"ahlerian structure $(\omega_{\rm can}, {^{\tiny{\mbox S}}}\!J, {^{\tiny{\mbox S}}}\!g)$ restricts to an almost K\"ahler structure on $(\Omega,g_{_f})$ and (2) in Theorem \ref{not exist} applies for $^{\tiny{\rm{S}}}\!J$ which is uniquely determined by $^{\tiny{\rm{S}}}\!g$ through $g$.
\begin{thm}\label{special case}  
Suppose that $M$ is a compact manifold. Let $L$ be a compact oriented exact Lagrangian submanifold of $T^*M$ relative to $\omega_{\rm can}$. Then the conclusions {\rm (1)} and ${\rm  (2)}$ in Theorem \ref{not exist} hold for the unique almost complex structure $^{\tiny{\rm S}}\!J$ compatible with $\omega_{\rm can}$ and the Sasakian metric $^{\tiny{\rm{S}}}\!g$ associated with a given Riemannian metric $g$ on $M$. 
\end{thm}

\section{Existence of holomorphic disks in K\"ahler surface}

Let $M$ be a K\"ahler manifold and let $D$ be the open unit disk in $\R^2$. The notion $u:(D,\p D)\to (M,L)$ denotes a mapping  $u:D\cup \p D\to M$ satisfying $u(\p D)\subset L$ where $L\subset M$ is a surface. 

Fix an orientation on $D$ and let $z=x+iy$ be the complex coordinate. This makes $D$ a Riemann surface. Let $\p D$ have the induced orientation from $D$. A smooth map $u$ as above pulls back the holomorphic tangent bundle of $M$ 
$$
T^{1,0}M=\big\{ X-iJX: X \in T_pM, p\in M\big\}$$
 to a smooth complex vector bundle 
$$
{\bf E}=u^*T^{1,0}M
$$ 
over $D$ of complex rank 2. The Riemannian metric $g$ on $M$ extends $\C$-linearly to $TM\otimes\C$. The map $u$ pulls it back to a complex bilinear form $\langle\cdot,\cdot\rangle$ on ${\bf E}$ and defines a Hermitian inner product $\langle\!\langle\cdot,\cdot\rangle\!\rangle$ on ${\bf E}$ via $\langle\!\langle U,V\rangle\!\rangle=\langle U,\overline{V}\rangle$ for any $U,V\in \bf E$. Since $J$ is parallel, the complex extension of the Levi-Civita connection of $(M,g)$ to $TM\otimes\C$ preserves $(1,0)$-type vectors in $TM\otimes\C=T^{1,0}M\oplus T^{0,1}M$, so $u$ pulls it back to a connection $\nabla$ on $\bf E$ which is compatible with $\langle\!\langle\cdot,\cdot\rangle\!\rangle$. The connection $\nabla$ decomposes into $\nabla=\nabla'+\nabla''$ where 
$$
\nabla': \A^{0,0}({\bf E}) \to \A^{1,0}({\bf E}),\,\,\,\,\nabla'' : \A^{0,0}({\bf E})\to\A^{0,1}({\bf E})
$$ 
and $\A^{p,q}({\bf E})$ is the space of ${\bf E}$-valued $(p,q)$-forms on $D$. In the complex coordinate $z$, $\A^{1,0}(\E)$ consists of multiples of $dz$ and $\A^{0,1}(\E)$ consists of multiples of $d\bar{z}$. 

The Koszul-Malgrange theorem (\cite{KM}, cf. \cite[Theorem 5.1]{AHS}) asserts that ${\bf E}\to D$ admits a unique holomorphic structure with respect to which $\nabla''$ is the Dolbeault $\overline\p$-operator on ${\bf E}$. In this holomorphic structure, a holomorphic section $W$ of ${\bf E}$ is characterized by any one of the following equivalent conditions
$$
\overline{\p} W=0\,\,\Longleftrightarrow\,\nabla'' W=0\,\,\Longleftrightarrow\nabla_{\frac{\p}{\p \bar z}}W=0
$$ 
where $\frac{\p}{\p \bar{z}} = \frac{1}{2}\big(\frac{\p}{\p x} +i\frac{\p}{\p y}\big)$. In terms of the real operators
$$
\overline\nabla_J=\frac{1}{2}\left(\nabla_{\frac{\p}{\p x}}+J\n_{\frac{\p}{\p y}}\right),\,\,\,\,\,\,\,
\nabla_J=\frac{1}{2}\left(\nabla_{\frac{\p}{\p x}}-J\n_{\frac{\p}{\p y}}\right).
$$
a section $V=X-iJX\in\Gamma(\bf E)$ is holomorphic if and only if $\overline\n_J X=0$;  and $W=Y-iJY\in\Gamma({\bf E})$ is anti-holomorphic if and only if $\n_J Y=0$.

The harmonic map equation can be written globally on $D$ as 
$$
\n_{\frac{\p}{\p\overline z}}\frac{\p u}{\p z}=0
$$
where $\frac{\p u}{\p z} =\frac{1}{2}\big( \frac{\p u}{\p x}-i\frac{\p u}{\p y}\big)$. When $u$ is harmonic, $\frac{\p u}{\p z}$ is a holomorphic section of ${\bf E}$, and it is straight to verify by using $\n J=0$ that $\overline{\p}_Ju-iJ\overline{\p}_Ju$ is an anti-holomorphic section of ${\bf E}$, where 
$$
\overline{\p}_J u = \frac{1}{2}\left(\frac{\p u}{\p x}+J\frac{\p u}{\p y}\right),\,\,\,\,\,\,\,
{\p}_J u = \frac{1}{2}\left(\frac{\p u}{\p x}-J\frac{\p u}{\p y}\right).
$$
A map $u:D\to M$ is holomorphic if $\overline{\p}_Ju =0$ (strictly speaking, $u$ is $(J_D,J)$-holomorphic where $J_D$ is the complex structure corresponds to $z$).
This setup is in \cite{ChenFraser2}, while a similar one for a general Riemannian manifold $M$ in terms of pullback of the full complexified tangent bundle $u^*TM\otimes\C$ appears in \cite{MM}. 

\vspace{.1cm}

Let $L$ be an oriented Lagrangian surface in $M$. Along the boundary $u(\p D)$ in the Lagrangian submanifold $L$ of $M$, $u$ pulls back $TL$ to a totally real vector bundle ${\bf F}\subset {\bf E}|_{\p D}$ over $\p D$ of real rank 2 with fibre at $z\in\p D$ given by 
$$
F_z =\left\{ X-iJX\, |\, X \in T_{u(z)}L\right\}.
$$
Take the orientations for $\E, {\bf F}$ via the pullback by $u$ (fibrewise). 

The boundary Maslov index $\mu({\bf E},{\bf F})$ of the bundle pair $({\bf E},{\bf F})$ can be formulated, as a Riemann-Roch theorem for the bundle pair $({\bf E},{\bf F})\to(D,\p D)$, in terms of the Fredholm index of a real Cauchy-Riemann operator and the Euler number of $D$, see \cite[Theorem C.3.5, Theorem C.1.10]{MS}. In differential geometric terms, $\mu({\bf E},{\bf F})$ admits an integral expression in the setting $u:(D,\p D)\to (M,L)$: 
\begin{equation*}
\mu_u(D, \p D) = \frac{i}{\pi} \int_D u^*\mbox{tr}(R) -\frac{1}{\pi}\int_{\p D} \xi_J
\end{equation*}
where the 1-form $\xi_J$ on $L$ is $\R$-valued and $i\xi_J$ is the connection 1-form of the unitary connection of the Hermitian metric $\langle\!\langle\cdot,\cdot\rangle\!\rangle$ on ${\bf E}$ and $R$ is its curvature. The integer $\mu_u(D,\p D)$ is the boundary Maslov index $\mu({\bf E},{\bf F})$ and it changes sign if the orientation of $D$ is reversed, see \cite{Pacini}. The above Chern-Weil type formula is established for totally real $L$ and general compact bordered surfaces in \cite{Pacini}. When $M$ is K\"ahler, 
\begin{equation}\label{Maslov}
\mu_u(D,\p D) = \frac{1}{\pi}\int_Du^*{\mbox{Ric}}_M -\frac{1}{\pi} \int_{\p D}u^*\omega(H,\cdot) 
\end{equation}
where $H$ is the mean curvature vector of $L$, $\omega(H,\cdot)$ is the interior product of the K\"ahler form $\omega$ with $H$ (see  \cite{Da}, \cite{Mor}, \cite{C-G}, \cite{Ono}). 

\vspace{.2cm}

\noindent{\bf Proof of Theorem \ref{represent}.} For each $\gamma\in \ker i_*$ and $\gamma$ is not contractible in $L$ to a single point, let 
$
u:(D,\p D)\to(M,L)
$
be an area minimizing branched immersion with free boundary such that $u(\p D)$ is homotopic to $\gamma$ on $L$. The free boundary condition means that $u(D)$ is orthogonal to $L$ along $\p D$. Such an area minimizing map is necessarily nonconstant as $\gamma$ is not contractible in $L$. The existence of $u$ is due to R. Ye \cite{Ye}, also see \cite[Theorem 1 (ii)]{ChenFraser1}. The map $u$ is a branched minimal immersion. Since $L$ is compact, $u$ is smooth on ${\overline D}=D\cup\p D$ (\cite[Proposition 1]{Ye}, cf. \cite{Jost1}, \cite[Theorem 2]{Gruter}). 

\vspace{.1cm}

{\bf Case 1.}  $\mu_u(D,\p D)\geq 0$. 

Define the space of admissible holomorphic sections of $\bf E$ by 
 $$
\mathcal H=\big\{V\in \Gamma({\bf E}):\overline{\partial} V =0 \mbox{ on $D$, Re\,$V(z)\in T_{u(z)}L$ on $\p D$}\big\}.
$$
According to \cite[Theorem 2.2]{ChenFraser2} (cf. \cite[Theorem C.1.10]{MS}) 
$$
\dim_{\R}{\mathcal H}\geq \mu_u(D,\p D) + 2\chi(D)\geq 2
$$  
where $\chi(D)$ is the Euler characteristic number of $D$ and $\mu_u(D,\p D)\geq 0$.  
Therefore, there exist two linearly independent elements $s_1,s_2\in\mathcal H$ over $\mathbb R$, {\it i.e.},  if for any real numbers $c_1,c_2$ such that $c_1s_1(z)+c_2s_2(z)=0$ at all $z$ then $c_1,c_2$ are zero. However, this does not prevent $f_1(z)s_1(z)+f_2(z)s_2(z)=0$ while $f_1(z),f_2(z)$ not all zero at each $z\in D$. 

\vspace{.1cm}

Since $u$ is harmonic, $\overline{\p}_Ju-iJ\overline{\p}_Ju\in\Gamma({\bf E})$ is anti-holomorphic. As in \cite{ChenFraser2}, to get an anti-holomorphic section section of ${\mathcal H}$, let
$$
W= (y I+xJ)(\overline{\p}_Ju-iJ\overline{\p}_Ju)\in\Gamma({\bf E})
$$
where $I$ is the identity map on $TM$ and $x+iy=z$ on $D$. We claim $W\in{\mathcal H}$. 
First of all, $W$ is anti-holomorphic: 
\begin{align*}
2\n_J \big[(y I +Jx)\overline{\p}_Ju\big]=\big( \n_{\frac{\p}{\p x}} -J\n_{\frac{\p}{\p y}}\big) \big[(y I+xJ)\overline{\p}_Ju\big]
= (y I+x J)\n_J\overline{\p}_Ju=0. 
\end{align*}
Second, ${\rm Re}\, W$ belongs to $TL$ along $u(\p D)$. To see this, observe that in the polar coordinates on $D$
$$
2\,{\rm Re}\,W= rJ\frac{\p u}{\p r} - \frac{\p u}{\p\theta}.
$$
Along $u(\p D)$,  $\frac{\p u}{\p \theta}$ is tangent to $L$ since $u(\p D)\subset L$ and $J\frac{\p u}{\p r}$ is tangent to $L$ due to the fact that $\frac{\p u}{\p r}$ is orthogonal to $L$ (as $u$ satisfies the free boundary condition) and $L$ is Lagrangian. Thus, ${\rm Re}\,W$ is tangent to $L$. So $W\in{\mathcal H}$. 

For $j=1,2$, the function
$$
f_j= \big\langle s_j,\overline{W} \big\rangle
$$
is holomorphic where $\overline{W}$ is the complex conjugate of $W$.  We claim that  $f_j$ takes real values along $\p D$. To see this, we write $s_j = V_j-iJV_j$ and use $\langle JV_j, {\rm Re}\, W\rangle|_{\p D}=0$ (since both $V_j$ and ${\rm Re}\, W$ are tangent to $L$ and $L$ is Lagrangian) to conclude
$$
f_j = \big\langle V_j-iJV_j, {\rm Re}\, W\big\rangle=2\big\langle Y_j, {\rm Re}\, W\big\rangle.
$$
Therefore $f_j$ is a constant which must be zero since $W(0)=0$. 

There are two cases: 
 
{\bf (A)} $s_1,s_2$ are linearly independent over $\R$ at some point $p$ in $\p D$. 

In this case, $u$ is holomorphic by \cite[Theorem 1.1 (1)]{ChenFraser2}. For completeness we present a proof. 

Since $s_1(p),s_2(p)$ are linearly independent over $\R$, they are nonzero vectors and $V_1(p),V_2(p)$ are linearly independent over $\R$. If $W(p)\not=0$, then $s_1(p),s_2(p)$ are both orthogonal to $W(p)$ as $\langle\!\langle s_j,W\rangle\!\rangle =f_j=0$, so they are linearly dependent over $\C$ because $\dim_\C\E_p=2$. For some numbers $a_j+ib_j\in\C$, not all zero for $j=1,2$, 
$$
(a_1+ib_1)s_1(p) + (a_2+ib_2)s_2(p)=0. 
$$
Equivalently,
$$
\sum_{j=1}^2 a_jV_j(p)+ J\sum^2_{j=1}b_jV_j=0.
$$
As $L$ is Lagrangian and $V_j(p)\in T_{u(p)}L$, it then follows 
$$
\sum_{j=1}^2 a_jV_j(p)=0= \sum^2_{j=1}b_jV_j(p).
$$
But this contradicts the linear independence of $\{s_1(p),s_2(p)\}$ over $\R$ since some $a_j$ or $b_j$ is nonzero. In conclusion, $W(p)=0$ if $s_1(p),s_2(p)$ are linearly independent over $\R$ for $p\in\p D$. The linear independence is an open condition, so there is an arc $\gamma\subset \p D$ containing $p$ such that $s_1,s_2$ are linearly independent over $\R$ at each point in $\gamma$. Then $W=0$ everywhere on $\gamma$. Take an anti-holomorphic local frame $\sigma_1,\sigma_2$ of $\E$ on a neighborhood $U$ in $D$ with $\gamma$ in the closure of $U$. Then $W=f^1\sigma_1+f^2\sigma_2$ for some anti-holomorphic functions $f^1,f^2$ on $U$. However $f^2,f^2$ must vanish on $U$ since they vanish along $\gamma$. This says $W\equiv 0$ on $U$ hence $W=0$ on 
$D$ by the unique continuation of holomorphic functions. This implies $\overline{\p}_Ju=0$, in other words, $u$ is holomorphic. 

\vspace{.1cm}

{\bf (B)} $s_1,s_2$ are linearly dependent over $\mathbb R$ at each point in $\p D$. 

In this case, for any $p\in\p D$ there exist two constants $a_1(p),a_2(p) \in\mathbb R$, not all zero, such that 
$$
a_1(p)s_1(p)+a_2(p)s_2(p)=0.
$$
Then the section of $\bf E$ defined by 
$$
X_p(z) = a_1(p)s_1(z)+a_2(p)s_2(z),\,\,\forall z\in D\cup\p D
$$
satisfies 
$X_p\in\mathcal H,X_p(p)=0$
and $X_p\not\equiv 0$ as $s_1,s_2$ are linearly independent over $\R$.

Since $X_p,{W}$ are in $\mathcal H$, $f_1=f_2=0$, it follows that for all $z\in D$ 
$$
\big\langle\!\big\langle X_p(z),{W}(z)\big\rangle\!\big\rangle= \big\langle X_p(z), \overline{W}(z)\big\rangle=a_1(p)f_1(z)+a_2(p)f_2(z)=0.
$$
This implies that 
$
X_p(z)\perp W(z)
$
with respect to the Hermitian metric $\langle\!\langle\cdot,\cdot\rangle\!\rangle$ on $\E_z$ whenever $W(z)\not=0$. By the same token, for any point $p_1\in \p D$, $p_1\not=p$, it also holds 
$
X_{p_1}(z)\perp W(z)
$
whenever $W(z)\not=0$. This means that both $X_p(z)$ and $X_{p_1}(z)$ are orthogonal to $W(z)$ at each $z$ where $W(z)\not=0$ (equivalently, $\frac{\p u}{\p \overline z}(z)\not=0$ and $z\not=0$). In particular, at such a point $z$, the vectors $X_p(z),X_{p_1}(z)$ are linearly dependent over $\C$ since $\dim_{\mathbb C}{\bf E}_{z}=2$. 

\vspace{.1cm}

{\bf B.1} $\frac{\p u}{\p \overline z}\equiv 0$ on ${D}$. This means $u$ is holomorphic. 

{\bf B.2} $\frac{\p u}{\p \overline z}\not\equiv 0$ on ${D}$. We shall show this cannot happen. 

Since $u$ is harmonic,  $\frac{\p u}{\p \overline z}$ is anti-holomorphic. Away from at most finitely many points, $\frac{\p u}{\p \overline z}\not=0$ on $D$, otherwise $u$ would be constant. 

Fix a point $p_1\in\p D$ and pick any $p\in\p D$ with $p\not=p_1$. Then $X_p(z)$ and $X_{p_1}(z)$ are linearly dependent over $\C$ at any point $z\in\overline{ D}\setminus\{0\}$ where $\frac{\p u}{\p \overline z}(z)\not=0$ since both of them are orthogonal to $W$ as saw before. The complement of the zero set of $\frac{\p u}{\p \overline z}$ is open and dense in $D$, by continuity, the vectors $X_p(z)$ and $X_{p_1}(z)$ are linearly dependent over $\C$ at any point $z\in\overline{D}$. For each pair $(z,p)\in\overline{D}\times\p D$, there exist $C_1(z,p,p_1),C_2(z,p,p_1)\in\C$ (we shall write $C_1,C_2$ for simplicity), not all zero, such that 
$$
C_1X_p(z) +C_2X_{p_1}(z)= 0.
$$
This means
\begin{equation}\label{C-linear}
C_1\big\{a_1(p)s_1(z)+a_2(p)s_2(z)\big\}+C_2\big\{a_1(p_1)s_1(z)+a_2(p_1)s_2(z)\big\}=0.
\end{equation}

We can express $s_1,s_2\in{\bf E}$ as 
$$
s_1=Y_1-iJY_1, \,\,\,\,s_2=Y_2-iJY_2
$$
where $Y_1,Y_2$ are sections of the real bundle $u^*TM$, and write 
$$ 
C_1=A_1+iB_1, \,\,\,\,C_2=A_2+iB_2
$$ 
where $A_1,A_2,B_1,B_2$ are $\R$-valued functions depending on $z,p,p_1$ (again, we omit the dependence on $p,z,p_1$ for notational simplicity). 

Regrouping terms in \eqref{C-linear}
\begin{equation}\label{C-linear:re}
\big\{C_1a_1(p)+C_2a_1(p_1)\big\}s_1(z)+\big\{C_1a_2(p)+C_2a_2(p_1)\big\}s_2(z)=0.
\end{equation}
This leads to 
\begin{align*}
\mbox{The first term in } &\eqref{C-linear:re} =\big\{(A_1+iB_1)a_1(p)+(A_2+iB_2)a_1(p_1) \big\}(Y_1-iJY_1)(z) \\
=& \big\{A_1a_1(p)+A_2a_1(p_1)\big\}Y_1(z) + \big\{ B_1a_1(p)+B_2a_1(p_1)\big\} JY_1(z)\\
+&\,i\big\{B_1a_1(p) +B_2a_1(p_1)\big\} Y_1(z) -i\big\{ A_1a_1(p)+A_2a_1(p_1)\big\} JY_1(z)\\
:=& \,Z_1 -iJZ_1.
\end{align*}
Similarly, the second term in \eqref{C-linear:re} can be expressed as $Z_2-iJZ_2$. Rewrite \eqref{C-linear}: 
\begin{equation}\label{C-linear'}
(Z_1+Z_2)-iJ(Z_1+Z_2)=0.
\end{equation}
Noting $Z_1,Z_2$ are real, we conclude 
\begin{equation}
Z_1+Z_2=0.
\end{equation}
In other words, 
\begin{align*}
&\big\{A_1a_1(p)+A_2a_1(p_1)\big\}Y_1(z) + \big\{ B_1a_1(p)+B_2a_1(p_1)\big\} JY_1(z)\label{1}\\
+&\big\{A_1a_2(p)+A_2a_2(p_1)\big\}Y_2(z) + \big\{ B_1a_2(p)+B_2a_2(p_1)\big\} JY_2(z)=0.
\end{align*}

Since $L$ is Lagrangian and $Y_1(z),Y_2(z)\in T_{u(z)}L$ for $z\in\p D$, we are led to 
\begin{eqnarray}
&&\big\{A_1a_1(p)+A_2a_1(p_1)\big\}Y_1(z)+ \big\{A_1a_2(p)+A_2a_2(p_1)\big\}Y_2(z) =0\label{1}\\
&&\big\{B_1a_1(p)+B_2a_1(p_1)\big\} JY_1(z)+ \big\{ B_1a_2(p)+B_2a_2(p_1)\big\} JY_2(z)=0.\label{equation2}
\end{eqnarray}
Letting $J$ acts on \eqref{equation2} we deduce
\begin{equation}\label{2}
\big\{ B_1a_1(p)+B_2a_1(p_1)\big\} Y_1(z)+ \big\{ B_1a_2(p)+B_2a_2(p_1)\big\} Y_2(z)=0.
\end{equation}

Noting $X_p(p)=0$, that is to say 
\begin{align*}\label{at p}
0&= a_1(p)\big\{Y_1(p)-iJY_1(p)\big\}+a_2(p)\big\{Y_2(p)-iJY_2(p)\big\}\\
&= \big\{a_1(p)Y_1(p)+a_2(p)Y_2(p)\big\} -i J\big\{a_1(p)Y_1(p)+a_2(p)Y_2(p)\big\}.\nonumber
\end{align*}
In turn
\begin{equation}\label{at p again}
a_1(p)Y_1(p)+a_2(p)Y_2(p)=0.
\end{equation}

Using \eqref{at p again} in \eqref{1} and \eqref{2} at $z=p$ shows 
\begin{align}
A_2\big\{a_1(p_1)Y_1(p)+a_2(p_1)Y_2(p)\big\}&=0\\
B_2\big\{a_1(p_1)Y_1(p)+a_2(p_1)Y_2(p)\big\}&=0
\end{align}
However, $A_2, B_2$ are not both zero. Then 
\begin{equation}
a_1(p_1)Y_1(p)+a_2(p_1)Y_2(p)=0.
\end{equation}
This implies 
\begin{align*}
X_{p_1}(p) &=\big\{a_1(p_1)Y_1(p)+a_2(p_1)Y_2(p)\big\} -iJ  \big\{a_1(p_1)Y_1(p)+a_2(p_1)Y_2(p)\big\}=0. \end{align*}

Since $p$ can be any point in $\p D$, we see $X_{p_1}\equiv 0$ along $\p D$. Then $X_{p_1}\equiv 0$ in $D$ by the holomorphicity of $X_p$. But this is impossible for $X_{p_1} = a_1(p_1)s_1+a_2(p_1)s_2$ with $s_1,s_2$ linearly independent on $D$ over $\mathbb R$ and $a_1(p_1),a_2(p_1)$ not both zero. This means that the case {\bf B.2}  cannot occur. 

Therefore  $u$ is holomorphic for {\bf Case 1}.

\vspace{.1cm}

{\bf Case 2.} $\mu_u(D,\p D)\leq 0$. 

Reverse the orientation on $D$ and take the complex coordinate $w = x - iy$. Denote the new Riemann surface by $\widetilde{D}$ which is $D$ with the complex structure defining $w$. Consider the bundle pair $(\E,{\bf F})$ now over $\widetilde{D}$ with the same bundle orientation (fiberwise). 
The same $\C$-linearly extended connection $\nabla$ on $\E$ (from the Levi-Civita connection on $M$) splits into $\widetilde{\nabla}' +\widetilde{\nabla}''$ as
$$
\widetilde{\nabla}': \widetilde{\A}^{0,0}(\E)\to\widetilde{\A}^{1,0}(\E),\,\,\,\,\,\widetilde{\nabla}'': \widetilde{\A}^{0,0}(\E)\to\widetilde{\A}^{0,1}(\E)
$$
where $\widetilde{\A}^{p,q}(\E)$ is the $(p,q)$-forms on $D$ with values in $\E$. In particular, $\widetilde{\A}^{1,0}(\E)$ consists of multiples of $dw$ and $\widetilde{\A}^{0,1}(\E)$ consists of multiples of $d\overline{w}$.

The Koszul-Malgrange theorem uniquely determines a holomorphic structure on the complex bundle ${\bf E}\to{\widetilde D}$. Denote its Dolbeault $\overline{\p}$-operator by $\widetilde{\p}$ (to distinguish the $\overline{\p}$-operator on ${\bf E}\to D$). Respect to this new holomorphic structure, a section $W$ of $\E$ is holomorphic if any one of the equivalent conditions below holds
 $$
 \widetilde{\p}W=0\Longleftrightarrow \,\,\widetilde{\nabla}''W=0 \,\,\Longleftrightarrow\,\,\nabla_{\frac{\p}{\p\overline w}}W=0
 $$ 
 where $\frac{\p}{\p \overline w}=\frac{1}{2}\big(\frac{\p}{\p x}-i\frac{\p}{\p y}\big)$. The space of admissible holomorphic variations is
$$
\widetilde{\mathcal H}=\left\{V\in \Gamma({\bf E}): \widetilde\p V=0 \mbox{ on $D$, Re $V(z)\in T_{u(z)}L$ on $\p D$}\right\}
$$
and Theorem 2.2 in \cite{ChenFraser2} says 
$$ 
\dim_{\mathbb R}\widetilde{\mathcal H}\geq 2+{\mu}_u(\widetilde{D},\p \widetilde{D}).
$$ 

For the fixed map $u$, when reversing the orientation of $D$, it is evident from the integral formula \eqref{Maslov} that the new Maslov index is the opposite of the original one, since the differential forms being integrated remain unchanged but the orientations on $D$ and $\p D$ are both reversed:
$$
{\mu}_u(\widetilde{D},\p \widetilde{D}) = -\mu_u(D,\p D)\geq 0.
$$
Therefore, 
$$
\dim_{\mathbb R}\widetilde{\mathcal H}\geq 2.
$$
The argument in {\bf Case 1} then goes through verbatim, so $u$ is holomorphic in $w$, hence anti-holomorphic in $z$. 

Finally, since the case $\mu_u(D,\p D)=0$ falls into both {\bf Case 1} and {\bf Case 2}, $u$ would be both holomorphic and anti-holomorphic for a fixed choice of orientation on $D$ and $z$, hence it must be constant. But this contradicts the nonconstancy of $u$. 
We conclude that $\mu_u(D,\p D)\not=0$. \hfill$\square$

\begin{rem}
The above proof of $\pm$ holomorphicity of $u$ uses, in an essential way, the fact that the radial derivative of $u$ is orthogonal to $L$ along the boundary of the free boundary minimal disk. There may well be other holomorphic disks with boundary in $L$ of zero boundary Maslov class representing the same relative homotopy class $\pi_1(M,L)$; however, these disks (if exist) cannot meet $L$ orthogonally. 
\end{rem}

\section{Exact Lagrangian embeddings in almost K\"ahler manifolds}

Holomorphic disks are useful to study the question about whether a compact surface is realizable as an exact Lagrangian surface in an exact symplectic 4-manifold, as demonstrated in \cite{Gromov}. A symplectic manifold $M$ is exact if its symplectic 2-form $\omega=d\lambda$ for some differential 1-form $\lambda$ on $M$; in such an exact $M$ a Lagrangian submanifold $L$ is exact if $\lambda|_L$ is an exact 1-form on $L$. For example, the cotangent bundle of a smooth manifold with the canonical symplectic structure is exact and the zero section is an exact Lagrangian submanifold. Exact symplectic manifolds are necessarily noncompact by Stokes' theorem since $\omega$ is exact and nondegenerate. 

It is tempting to explore where Theorem \ref{represent} may lead to in this direction.

\subsection{Homotopy Lemma with Lagrangian constaint} 
Let $\vp: N\to M$ be a smooth map between almost Hermitian manifolds, one can define, from the complexified differential $d\vp:TN\otimes\C\to TM\otimes\C$, the partial differentials 
\begin{align*}
&\p\vp: T^{1,0}N \to T^{1,0}M\\
&\overline{\p}\vp: T^{0,1}N\to T^{1,0}M\\
&\p\overline{\vp}:T^{1,0}N\to T^{0,1}M\\
&\overline{\p}\overline{\vp}:T^{0,1}N\to T^{0,1}M
\end{align*}
and the partial energy densities 
\begin{align*}
e'(\vp)&= |\p\vp|^2,\,\,\,\,e''(\vp)= |\overline{\p}\vp|^2. 
\end{align*}
The energy density of $\vp$ is 
$$
e(\vp) = e'(\vp)+e''(\vp).
$$ 
Denote the energies by 
$$
E(\vp)=\int_N e(\vp)dv_g,\,\,\,E'(\vp)=\int_N e'(\vp)dv_g,\,\,\,\,E''(\vp)=\int_N e''(\vp)dv_g.
$$

Lichnerowicz's homotopy lemma in \cite{Li} admits a natural extension to almost K\"ahler manifolds with boundary which is mapped to a Lagrangian submanifold of the (target) almost K\"ahler manifold: 

\begin{lem}\label{homotopy} 
Let $L$ be a Lagrangian submanifold in an almost K\"ahler manifold $M$ and let $N$ an almost K\"ahler manifold such that $N\cup\p N$ is a compact manifold with boundary $\p N$. Let $\vp_t: (N,\p N)\to (M,L)$ for $t\in[0,1]$ be a smooth family of maps. Then 
\begin{enumerate}
\item[(i)] $E'(\vp_t) - E''(\vp_t)$ is a constant which depends on the homotopy class of $\vp_0$. 
\item[(ii)] If $\omega_{_M}=d\lambda$ where $\lambda$ is a ${\rm 1}$-form on $M$ and is exact on $L$, then $E'(\vp_t)=E''(\vp_t)= \frac{1}{2}E(\vp_t)$. 
\end{enumerate}
\end{lem} 
\begin{proof} An observation of Lichnerowicz in \cite{Li} (cf. \cite[Homotopy Lemma (8.8)]{E-L}) for smooth maps between almost K\"ahler manifolds asserts
$$
e'(\vp_t)-e''(\vp_t) = \big\langle \omega_{_N},  \vp_t^*\omega_{_M}\big\rangle
$$
and 
$$
\frac{\p}{\p t}\big(\vp_t^*\omega_{_M}\big) = d\left(\vp^*_t\omega_{_M}\big(\frac{\p \vp_t}{\p t}, \cdot\big)\right).
$$
By Stokes' theorem 
\begin{align*}
\frac{d}{dt}\int_N \big(e'(\vp_t)-e''(\vp_t)\big)dv_g&=\frac{d}{dt}\int_N  \big\langle \omega_{_N}, \vp_t^*\omega_{_M}\big\rangle dv_g\\
&=\int_N \big\langle\omega_{_N}, \frac{\p}{\p t}\big(\vp^*_t\omega_{_M}\big)\big\rangle dv_g\\
&= \int_{N} d\left(\vp^*_t\omega_{_M}\big(\frac{\p \vp_t}{\p t}, \cdot\big)\right)\wedge *_g \omega_{_N}\\
&= \int_{N} d\left(\vp^*_t\omega_{_M}\big(\frac{\p \vp_t}{\p t}, \cdot\big)\wedge *_g \omega_{_N}\right)\\
&=\int_{\p N} \left(\vp^*_t\omega_{_M}\big(\frac{\p \vp_t}{\p t}, \cdot\big)\right)\wedge *_g\omega_{_N}
\end{align*}
where $*_g$ denotes the Hodge star operator of the Riemannian manifold $(N,g)$ and 
$$
d(*_g\omega_{_N}) = \frac{1}{(n-1)!}d\omega_{_N}^{n-1}=0.
$$
where $n=\dim_\C N$. 

Since $\vp_t(\p D)\subset L$ for all $t$, we have $\frac{\p\vp_t}{\p t}(x)\in T_{\vp_t(x)}L$ for any $x\in\p N$. Then, it follows from the fact that $L$ is Lagrangian
$$
\vp^*_t\omega_{_M}\big(\frac{\p \vp_t}{\p t}, X\big)= \omega_{_M}\big(\frac{\p \vp_t}{\p t}, \vp_*X\big)=0
$$ 
for any $X\in T_xN$. We conclude $E'(\vp_t)-E''(\vp_t)$ is a homotopy invariant. 

\vspace{.1cm}

When $\omega_{_M}=d\lambda$ on $M$, 
\begin{align*}
e'(\vp_t)-e''(\vp_t)&= \big\langle\omega_{_N},\vp_t^*\omega_{_M} \big\rangle\\
&=\vp_t^*\omega_{_M}\wedge *_g\omega_{_N}\\
&= d\big(\vp_t^*\lambda\big)\wedge *_g\omega_{_N}\\
&=d\left(\vp_t^*\lambda\wedge *_g\omega_{_N} \right).
\end{align*}
Restricting along $L$, we know $\lambda= df$ for some smooth function $f$ on $L$ by assumption. Thus 
$$
\vp_t^*\lambda\wedge *_g\omega_{_N} = d\big( f\circ \vp_t\big)\wedge *_g\omega_{_N}=d\big(f\circ\vp_t \,\omega_{_N}^{n-1}\big).
$$
This implies, via Stokes' theorem, 
\begin{equation}
E'(\vp_t)-E''(\vp_t)=\int_{\p N} d\big(f\circ\vp_t \,\omega_{_N}^{n-1}\big)=0.
\end{equation}
In particular,
$$
\frac{1}{2}E(\vp_t) =E'(\vp_t)=E''(\vp_t).
$$
\end{proof}

\noindent{\bf Proof of Theorem \ref{no exact}.}
Since there is a loop $\gamma\subset L$ which is nontrivial in $\pi_1(L)$ but bounds a topological disk in $M$, by Theorem \ref{represent} there exists a nonconstant $J$-holomorphic map $u:(D,\p D)\to (M,L)$. From Lemma \ref{homotopy} (with $N=D$), this is impossible when $L$ is exact since $E(u)$ would be 0.
\hfill$\square$

\vspace{.1cm}

\noindent{\bf Proof of Theorem \ref{not exist}.} (i) By Lemma \ref{homotopy} (with $N=D$), any $J$-holomorphic map $u:(D,\p D)\to (M,L)$ must be constant since $E''(u)=0$. (ii) The existence of the map $u$ comes from the existence theory in \cite{Ye} (cf. \cite{ChenFraser1}) while its minimizing properties for $E',E''$ follows from Lemma \ref{homotopy}. \hfill$\square$

\subsection{Cotangent bundles} 

Suppose that $M$ is an $n$-dimensional manifold. Let $x=(x^1,...,x^n):U\to\mathbb R^n$ be a system of local coordinates in an open set $U\subset M$. The cotangent bundle $\pi: T^*M\to M$ has a local chart $(x^1,...,x^n, y_1,..., y_n)$ on $\pi^{-1}(U)$ where $y_i$'s are the components of covector fields $v^*\in T^*U$ written as $v^* = y_i dx^i.$ The coordinates system $T^*U\to \mathbb{R}^n\times\mathbb{R}^n$ is given by $(q,v^*)\to( x\circ\pi(q), y(q,v^*))$. The canonical symplectic structure on $T^*M$ is $\omega_{\footnotesize{\mbox{can}}}=-d\lambda$ where $\lambda= y_idx^i$ is the canonical 1-form, and $\omega_{\footnotesize{\mbox{can}}}$ is known to be be independent of the choice of the local coordinates. 

Let $g$ be a Riemannian metric on $M$. A basis of $T(T^*U)$ can be chosen as 
$$
{\mathscr B}=\left\{X_1,...,X_n, \frac{\p}{\p y_1},...,\frac{\p}{\p y_n}\right\}
$$ 
where (summation over repeated indices is assumed) 
$$
X_i = \frac{\p}{\p x^i}+\Gamma^k_{is} y_{_k}\frac{\p}{\p y_s}
$$
and $\Gamma^k_{ij}$ are the Christoffel symbols of the Levi-Civita connection of $g$ in the local coordinates $(x^1,...,x^n)$. The bundle projection $\pi|_{_U}:T^*U\to U$ satisfies 
$$
d\pi|_{_U}(X_i) = \frac{\p}{\p x^i}, \,\,\,\,\,d\pi|_{_U}\big(\frac{\p}{\p y_i}\big)=0.
$$

The Sasakian metric $^{\tiny{\mbox S}}\!g$ on $T^*M$ can be defined (cf. \cite{SaAg}) in the basis $\mathscr B$ by 
\begin{equation}\label{metric1}
\left\{
\begin{aligned}
&\,\,^{\tiny{\mbox S}}\!g\big(X_i,X_j\big) =  g_{ij}\\
&\,\,^{\tiny{\mbox S}}\!g\big(X_i,\frac{\p}{\p y_j}\big)= 0\\
&\,\,^{\tiny{\mbox S}}\!g\big(\frac{\p}{\p y_i},\frac{\p}{\p y_j}\big) = g^{ij}. 
\end{aligned}\right.
\end{equation}

\begin{prop}\label{complete}
If $(M,g)$ is a compact Riemannian manifold, then the Sasakian metric $^{\tiny{\rm S}}\!g$ on $T^*M$ is complete. 
\end{prop}
\noindent{\bf Proof.} Since $M$ is compact, there exists a finite cover $\big\{U_1,...,U_m\big\}$ of $M$ by normal coordinate ball $U_i$'s such that on each of them, the metric $g$ is uniformly equivalent to the Euclidean metric:
\begin{equation}\label{metric}
\frac{1}{2}  \sum_i dx^i\otimes dx^i\leq  g\leq 2 \sum_idx^i\otimes dx^i.
\end{equation}
Any tangent vector $X$ of $T^*M$ can be written as $X=a^i X_i +b_j\frac{\p}{\p y_j}$ in the basis $\mathscr B$ and its length can be estimated
\begin{equation}\label{length}
\|X\|^2_{^{\tiny\mbox S}\!g} = \sum_{i,j}g_{ij}a^ia^j +\sum_{i,j}g^{ij}b_ib_j\geq \frac{1}{2}|a|^2+\frac{1}{2}|b|^2
\end{equation}
where $|\cdot |$ denotes the Euclidean length of vectors. 

Let $d_{^{\tiny{\mbox S}}\!g}, d_g$ be the distance functions and let $^{\tiny{\mbox S}}\nabla,\nabla$ be the Levi-Civita connections of $(T^*M,{^{\tiny{\mbox S}}\!g}),(M,g)$, respectively. It follows from \cite[Theorem 2.1, i), iv)]{SaAg} that every fibre $\pi^{-1}(x)$ is totally geodesic and so is the zero section of $T^*M$ (identified with the inclusion of $M$ in $T^*M$). 

Using the total geodesicity of the fibres and the zero section together with the fact that the induced metric on $\pi^{-1}(x)$ from $^{\tiny{\mbox S}}\!g$ is $g^{-1}(x)$, we have 
\begin{eqnarray*}
d_{^{\tiny{\mbox S}}\!g}\big((p,0), (x,y)\big)&\geq& d_{^{\tiny{\mbox S}}\! g}\big((x,y),(x,0)) - d_{^{\tiny{\mbox S}}\! g}\big((p,0),(x,0)\big)\\
&=& \ell_{^{\tiny{\mbox S}}\!g}\big(\big\{(x,tp+(1-t)y): 0\leq t\leq 1\big\}\big) -d_g\big((p,0),(x,0)\big)\\
&\geq& C(g)|y| -C(g)(|x|+1).
\end{eqnarray*}
Consequently 
$$
d_{^{\tiny{\mbox S}}\!g}((x,y),(p,0))\to+\infty\,\,\,\mbox{ as $|y|\to+\infty$.} 
$$

We now construct an exhaustion of $T^*\Sigma$ by a sequence of relatively compact open sets as follows. Let 
$$
T^*M =  \bigcup_{j=1}^\infty\bigcup_{\alpha=1}^m{ K^\alpha_j}
$$
where 
$$
K^\alpha_j= \big\{ (x,y): x\in U_\alpha, y\in\pi^{-1}(x), |y|< j\in\mathbb N\big\}.
$$
It is clear
$$
\bigcup^m_{\alpha=1}K^\alpha_j \subset \bigcup^m_{\alpha=1}K^\alpha_{j+1} \subset\subset T^*M.
$$
If $q_n\not\in \bigcup^m_{\alpha=1}K_n^\alpha$ then the sequence $\{q_n\}$ can be written as a union of at most $m$ subsequences and each of them belongs to $\pi^{-1}(U_\alpha)$ for some $\alpha\in\{1,..., m\}$. The previous analysis on a single chart then asserts that for each such subsequence (still denote by $\{q_n\}$ for simplicity) $d_{^{\tiny{\mbox S}}\!g}(q_n,p)\to+\infty$ for any fixed $p\in M$. 
Hence $^{\tiny{\mbox S}}\!g$ is complete by the Hopf-Rinow theorem. 
\hfill$\square$

\vspace{.1cm}

The Sasakian metric $^{\tiny{\mbox S}}\!g$ and $\omega_{\rm can}$ uniquely determine an almost complex structure $^{\tiny{\mbox S}}\!J$ on $T^*M$ that satisfies: at each $p\in T^*M$, in the local basis ${\mathcal B}$, 
\begin{equation}\label{J}
^{\tiny{\mbox S}}\!J(X_i) = g_{ij}\frac{\p}{\p y_j},\,\,\,\, ^{\tiny{\mbox S}}\!J\frac{\p}{\p y_i} = - g^{ij}X_j
\end{equation}
and 
\begin{equation}\label{sym form}
\omega_{\rm can}(V,W)(p) = {^{\tiny{\mbox S}}\!g}({^{\tiny{\mbox{S}}}\!J}V,W)(p)
\end{equation}
for any $V,W\in T_pT^*M$. The triple $\{\omega_{\rm can}, {^{\tiny{\mbox S}}\!g}, {^{\tiny{\mbox S}}\!J}\}$ then defines an almost K\"ahlerian structure on $T^*M$.  

\vspace{.1cm}

It is shown in \cite[(2.12)]{SaAg} that the connection $^{\tiny{\mbox S}}\nabla$ on $T^*M$, when expressed in the basis ${\mathscr B}$ as 
$$
^{\tiny{\mbox S}}\nabla_{v_\alpha}v_{\beta} = {^{\tiny{\mbox S}}}\Gamma^\gamma_{\alpha\beta}v_{\gamma},
$$
with $\alpha,\beta,\gamma=1, ... , 2n$, satisfies 
\begin{equation}\label{connection}
^{\tiny{\mbox S}}\Gamma= C(g) * \Gamma (g)+C(g)*y
\end{equation}
where (and in sequel) the notion $C(g)$ denotes uniform constants (may involve indices from coordinates) only depends on $(M,g)$ and the expression $A*B$ refers to multiplications and is linear in $A,B$. 
The $(4,0)$-type Riemannian curvature tensor $^{\tiny{\mbox S}}\!R$ of $^{\tiny{\mbox S}}\!g$ has the following form in the basis ${\mathscr B}$ (see \cite[(3.1)]{SaAg}):
\begin{equation}\label{R^S}
^{\tiny{\mbox S}}\!R= C(g) +C(g) * y +C(g)*y*y.
\end{equation}

In general, the sectional curvatures of $^{\tiny{\mbox S}}\!g$ are unbounded even when $M$ is compact, unless $(M,g)$ is flat. Nonetheless, it is evident from \eqref{R^S} that for compact $M$ the sectional curvatures of $^{\tiny{\mbox S}}\!g$ blow up at most quadratically in $|y|$ as $|y|\to+\infty$. 

\medskip

\noindent{\bf Proof of Theorem \ref{special case}}. 
Take a Riemannian metric $g$ on $M$. The cotangent bundle $T^*M$ is complete for the Sasakian metric $^{\tiny{\mbox S}}\!g$, by Proposition \ref{complete}. Since $L$ is compact, there exists a relatively compact open set $\Omega\subset T^*M$, in the topology determined by $d_{^{\tiny{\mbox S}}\!g}$, containing both  $L$ and $M$ (identified with the zero-section of $T^*M$). Let $\big\{U_1,...,U_m\big\}$ be the finite cover of $M$ by normal balls with \eqref{metric} as before. Define a ``doubling" of $\Omega$ by 
$$
\Omega'=\left\{(x,y): \big(x, \frac{y}{2}\big)\in\Omega\cap \pi^{-1}(U_\alpha) \mbox{ for some $\alpha=1,... ,m$}\right\}.
$$

Let $\{\eta_\alpha\}$ be a partition of unity subordinate to the covering $\big\{U_\alpha:\alpha=1,...,m\big\}$ of $M$. Define a smooth nonnegative function $f$ on $T^*M$: 
$$
f(x,y)=\left\{\begin{array}{rrr}
0, \hspace{1.8cm} &\mbox{for $(x,y)\in \Omega$}\\
\displaystyle\sum^m_{\alpha=1}\eta_\alpha(x)|y|^2, & \mbox{for $(x,y)\in \pi^{-1}(U_\alpha)\cap T^*\Sigma\setminus \Omega'$}
\end{array}\right. 
$$
and extend $f$ smoothly in $\Omega'\setminus\Omega$ as a nonnegative function with 
$$
|\p f|, |\p^2f|<C \big(1+|y|^2\big)
$$
for some uniform constant $C$ depends on $m,g$ where $\p$ denotes derivative in the coordinates of $\pi^{-1}(U_\alpha)$. We can further assume,  for any $(x,y)\in T^*\Sigma$
$$
 \big(1+|y|^2\big)e^{-f(x,y)}\leq C(g,\Omega).
 $$
As $\Omega$ is fixed, we shall write $C(g)$ for uniform constants as before in the sequel. 

Define a Riemannian metric $g_{_f}= e^{2f} \,{^{\tiny{\mbox S}}\!g}$ on $T^*M$. Since $^{\tiny{\mbox S}}\!g$ is complete and $e^{2f}\geq 1$, the metric $g_{_f}$ is complete. Under the conformal change of metrics, the $(4,0)$-type Riemannian curvature tensor transforms according to the well-known formula
\begin{equation}
R_{g_{_f}} = e^{2f}\left( ^{\tiny{\mbox S}}\!R - {^{\tiny{\mbox S}}\!g} \owedge\left({\rm{Hess}}_{\,^{\rm{S}}\!g}f-df\circ df+\frac{1}{2}|df|_{^{\rm S}\!g}^2\,{^{\tiny{\mbox S}}\!g}\right)\right):=e^{2f}L(f, {^{\tiny{\mbox S}}\!g})
\end{equation} 
where $\owedge$ is the Kulkarni-Nomizu product. 
By \eqref{connection} and \eqref{R^S}, $L(f,{^{\tiny{\mbox S}}\!g})$ has polynomial growth in $|y|$ as $|y|\to+\infty$, in fact 
\begin{equation}\label{L}
|L(f,{^{\tiny{\mbox S}}\!g})| \leq C(g) \big(1+ |y|^4\big).
\end{equation}

{\it Claim.}  $(T^*M,g_{_f})$ is homogeneously regular. 
\vspace{.1cm}

This will be verified in three steps. 

\vspace{.1cm}

{\bf Step 1.} Boundedness of sectional curvatures. 

\vspace{.1cm}

To estimate the sectional curvature $K_{g_{_f}}(\sigma)$ of $g_{_f}$ at any $p\in T^*M$ for an arbitrary 2-plane $\sigma$ in $T_pT^*M$, it is convenient to represent $\sigma$ by two $g_{_f}$-orthogonal nonzero vectors $X,Y$. Then 
\begin{eqnarray}
\big|K_{g_{_f}}(\sigma)\big|&=&\frac{\big|R_{g_{_f}}(X,Y,X,Y)\big|}{\|X\|^2_{g_{_f}}\|Y\|^2_{g_{_f}}}\\
&=&\frac{e^{2f}\big|L(f,{^{\tiny{\mbox S}}\!g})\big| |X|^2|Y|^2}{e^{4f}\|X\|^2_{^{\tiny{\mbox S}}\!g}\|Y\|^2_{^{\tiny{\mbox S}}\!g}}\nonumber\\
&\leq& 2e^{-2f} C(g)\big(1+|y|^4\big)\nonumber
\end{eqnarray}
by \eqref{length}, \eqref{L}. We conclude the boundedness of the sectional curvature of $g_{_f}$:
\begin{equation}\label{K}
\big|K_{g_{_f}}(\sigma)\big| \leq C(g). 
\end{equation}

{\bf Step 2.} Noncollapsing of volume of geodesic balls.

\vspace{.1cm}

For any $p\in T^*M$, let $U_p$ be a normal neighbourhood centred at $\pi(p)$ in $M$ satisfying \eqref{metric}. In the sequel , summing over repeated indices is understood. For any $(x,y)\in \pi^{-1}(U_p)$ and $W\in T_{(x,y)}T^*M$, in the basis ${\mathscr B}$ 
\begin{align*}
W&= a^iX_i+b_i\frac{\p}{\p y_i},\\
\|W\|_{g_{_f}}^2&=e^{2f} \left( g_{ij}a^ia^j+g^{ij}b_ib_j\right)\geq \frac{1}{2}e^{2f}\big(|a|^2+|b|^2\big).
\end{align*}
Measuring in the Euclidean metric:
\begin{equation}\label{euc}
\big\langle \frac{\p}{\p x^i},\frac{\p}{\p x^j}\big\rangle_{\rm euc}=\delta_{ij}=\big\langle \frac{\p}{\p y_i},\frac{\p}{\p y_j}\big\rangle_{\rm euc}, \,\,\,\,\big\langle \frac{\p}{\p x^i},\frac{\p}{\p y_j}\big\rangle_{\rm euc}=0.
\end{equation}
Recalling $X_i=\frac{\p}{\p x^i}+\Gamma^k_{il}y_k\frac{\p}{\p y_{_l}}$, then 
\begin{align*}
\|W\|_{\rm euc}^2&= a^ia^j\big\langle X_i,X_j\big\rangle_{\rm euc} + 2a^ib_j\big\langle X_i,\frac{\p}{\p y_j}\big\rangle_{\rm euc}+b_ib_j\big\langle \frac{\p}{\p y_i},\frac{\p}{\p y_j}\big\rangle_{\rm euc}\\
&= |a|^2+ a^ia^j\Gamma^k_{iq} \Gamma^l_{jq} y_k y_l  +2a^ib_j\Gamma^k_{ij}y_k+|b|^2\nonumber\\
&\leq C(g) e^{2f} \big(|a|^2+|b|^2\big).
\end{align*}
Therefore 
\begin{equation}\label{length1}
\|W\|_{\rm euc}^2\leq C(g)\|W\|_{g_{_f}}^2.
\end{equation}
It follows that the eigenvalues of $g_{_f}$ (as matrix in its coordinate representation) are no less than $C(g)^{-1/2}$; in particular, the volume element has a uniform lower bound
\begin{equation}\label{vol}
\sqrt{\det g_{_f}}\, dx^1dx^2dy_1dy_2\geq C(g) dx^1dx^2dy_1dy_2
\end{equation}
for some uniform constant $C(g)$. 

Another implication of \eqref{length1} is: for any smooth curve $\gamma$ in $\pi^{-1}(U_p)$ from $p$ to $(x,y)$ parametrized by $s$ 
\begin{align*}
\ell_{g_{_f}}(\gamma)&=\int_p^{(x,y)}  \|\gamma'(s)\|_{g_{_f}}ds\geq C(g)\int^{(x,y)}_p\|\gamma'(s)\|_{\rm euc}ds
\geq C(g)\,\ell_{\footnotesize{\mbox{euc}}}(\gamma)
\end{align*}
for some positive constant $C(g)$ which we write as $c(g)^{-1}$. Then the ball $B_p(r)\subset \pi^{-1}(U_p)$ in the metric $g_{_f}$ contains the ball ${\mathbb B}_p(c(g)r)$ measured in the Euclidean metric. Then, 
\begin{equation}\label{v}
\vol_{g_{_f}}\big(B_p(r)\big)\geq \vol_{g_{_f}}\big({\mathbb B}_p\big(c(g)r\big)\big) \geq C(g)\vol_{\footnotesize{\mbox{euc}}}\big({\mathbb B}_p\big(c(g)r\big)\big)\geq v_0>0.
\end{equation}
for $c(g)r\geq r_0$ where $r_0$ is the radius of the normal geodesic ball $U_p$ in $M$ centred at $\pi(p)\in M$ on which \eqref{metric} holds, and we used \eqref{vol} in the second inequality. 

Since $M$ is compact, it is easy to see \eqref{v} holds for any point $p$ in $T^*M$. 

\vspace{.1cm}

{\bf Step 3.} Uniform lower bound on injectivity radius.

\vspace{.1cm}

With \eqref{K} and \eqref{v} established, there is a positive lower bound on the injectivity radius of $(T^*M,g_{_f})$ according to \cite{CGT}. Then $(T^*M,g_{_f})$ is homogeneously regular, as claimed. 

\vspace{.1cm}

In the homogeneously regular space $(T^*M,g_{_f})$, there exists a nonconstant area minimizing disk $u(D)$ with free boundary $u(\p D)$ in $L$ representing a nontrivial $[\gamma]$ in $\pi_1(L)$. We now show the disk lies inside a prechosen region.  

The complete manifold $(T^*M, g_{_f})$ can be isometrically embedded into some Euclidean space $\R^N$ by Nash's isometric embedding theorem. The second fundamental form $A_\varphi$ of this embedding $\varphi:T^*M\to\R^N$ is bounded. Indeed, under the conformal change $g_{_f}=e^{2f} {^{\tiny{\mbox S}}}\!g$, the Levi-Civita connections of $g_{_f}$ and $^{\tiny{\mbox S}}\!g$ are related by 
\begin{equation}\label{A}
^f\nabla_XY= {^{\tiny{\mbox S}}}\nabla_XY+df(X)Y+df(Y)X-{^{\tiny{\mbox S}}}\!g(X,Y) \,{\rm grad}_{\,^{\tiny{\mbox S}}\!g}f
\end{equation}
where the unboundedness arises in $y$ of  a polynomial form by \eqref{connection} and \eqref{metric1}. 
From Gauss's formula 
$$
A_\varphi(X,Y)={^{\R^N}}\!\nabla_XY-{^f}\nabla_XY
$$
there is a constant $C_1(g)$ independent of $\Omega$ such that 
\begin{equation}\label{A-bound}
|A_\varphi|^2_{g_{_f}}=e^{-2f} |A_\varphi |^2_{^{\tiny{\mbox{S}}}\!g}\leq C_1(g).
\end{equation}

For the composition $D\overset{u}{\to} T^*M\overset{\varphi}{\to} \R^N$, the mean curvatures are related as
\begin{equation}\label{mean}
H_{\varphi\circ u}= du(H_u)+\nabla d{\varphi}(du, du)=\nabla d{\varphi}(du, du)
\end{equation} 
where in the last step we used the harmonicity of $u$. Thus $H_{\varphi\circ u}$ is bounded in $g_{_f}$ as $A_\varphi$ is by \eqref{A-bound}:
\begin{equation}\label{H}
|H_{\varphi\circ u}|\leq C_2(g).
\end{equation}

Although $\varphi$ and $N$ depend on $f$ hence on $\Omega$, the constants $C_1(g),C_2(g)$ do not. In other words, if we repeat the construction of $\Omega,f$ for a different $\Omega$, we can make the constants only depend on $g$. 

To use the discussion above, we first choose a topological disk $B$ in $T^*M$ with $\p B$ representing $[\gamma]$ in $\pi_1(L)$ to begin with, then choose $\Omega$ containing $B$ and get (from the existence theory for free minimal disks) the $g_{_f}$-minimal disk $u:(D,\p D)\to(T^*M,L)$. Being an area minimizer 
\begin{equation}\label{area}
{\rm Area}_{g_f}(u(D))\leq {\rm Area}_{g_f}(B)= {\rm Area}_{\,^{\tiny{\mbox S}}\!g}(B):= A_0
\end{equation}
since $g_{_f}={^{\tiny{\mbox S}}}\!g$ inside $\Omega$.

We claim that $\Omega$ can be chosen such that $u(D)$ stay entirely in $\Omega$.
According to \cite[Theorem 1, Theorem 2]{Ye}, 
$
|u|_{C^1}(\overline{D})\leq C(A_0).
$
This implies 
$$
\ell_{g_{_f}}(u(\p D))\leq \int^{2\pi}_0 \left| \frac{\p u}{\p \theta}\right| d\theta \leq 2\pi\,{C(A_0)}. 
$$
For any larger region containing $\Omega$ with the corresponding conformal metric constructed as before, the new conformal metric  is still $^{\tiny\mbox S}\!g$ inside $\Omega$. In particular, \eqref{area} holds with the same $A_0$ for the new minimizing disk. Therefore, the boundary length of the minimizing disk is bounded above by a constant only depends on $\Omega$. 

The estimate on the intrinsic diameter of a bordered surface in $\R^N$ \cite[Theorem 1.1]{Miu} says
\begin{equation}\label{diam}
{\rm{diam}}_{g_{_f}}(u(D))\leq C(N)\left(\int_{u(D)}|H_{\varphi\circ u}|_{g_{_f}} + \ell_{g_{_f}}(u(\p D)) \right).
\end{equation}
Then \eqref{H}, \eqref{area}, \eqref{diam} yield
\begin{equation}
{\rm{diam}}_{g_{_f}}(u(D))\leq C_3(g).
\end{equation}

Therefore, we can (initially) choose $\Omega$ sufficiently large such that $u(D)$ is contained in $\Omega$. Now $u(D)$ is a minimal surface in $(T^*M,{^{\tiny{\mbox S}}}\!g)$ with free boundary on $L$. The triple $(\omega_{\rm can}, {^{\tiny{\mbox S}}}\!J, {^{\tiny{\mbox S}}}\!g)$ restricts to an almost K\"ahler structure on $\Omega$. Theorem \ref{not exist} now applies to conclude Theorem \ref{special case} for the almost complex structure ${^{\tiny{\mbox S}}}\!J$.
\hfill$\square$

\bibliographystyle{alpha}

\end{document}